OpenStudy (anonymous):

Simplify: -7y - (2y - 9)

OpenStudy (anonymous):

-9(y+1)

OpenStudy (anonymous):

typo

OpenStudy (anonymous):

HUH!

OpenStudy (anonymous):

I hate typos

OpenStudy (anonymous):

Thank you, Gods of Algebra! I tip my hat to you! :)

OpenStudy (amistre64):

-7y -2y +9 = -9y+9 = -9(y+1) :) better

OpenStudy (anonymous):

I guess I'm confusing myself... it looks like you added the -7y and 2y... wouldn't that then be -5?

OpenStudy (amistre64):

first distribute the - into the paranthesis

OpenStudy (amistre64):

-(2y-9) = -2y +9

OpenStudy (amistre64):

now we add in the -7y -7y-2y +9 = (-7-2)y +9 = -9y+9